OpenStudy (anonymous):

Can you help me solve these polynomials?: y(squared) + 20y + 96

OpenStudy (anonymous):

=0?

OpenStudy (anonymous):

Or are you factoring?

OpenStudy (anonymous):

Yes factoring polynomials

OpenStudy (anonymous):

Ok, well set up your (y ) (y )

OpenStudy (anonymous):

(y+8)(y+12)

OpenStudy (anonymous):

Seriously dude? She wanted help not an answer

OpenStudy (anonymous):

i see :P

OpenStudy (anonymous):

divide 20 into 2 parts such that when multiplied, the product will be 96

OpenStudy (anonymous):

those are 12 and 8

OpenStudy (anonymous):

any way, notice that the two numbers multiply to equal c and add to equal b in the form ax^2 +bx +c

OpenStudy (anonymous):

another method: complete the squares

OpenStudy (anonymous):

y^2+20y+96=y^2+20y+100-4=(y+10)^2-2^2=(y+10+2)(y+10-2)=(y+12)(y+8)

OpenStudy (anonymous):

I do want help & answers here, my problem is knowing what's needed at which time. Here's another problem: 2x(cubed)y - x(squared)y + 5 xy(squared) +xy(cubed). I feel I should group here & my common factor is x(squared)y am I on the rt track with that?

OpenStudy (anonymous):

Will my answer end up being (2x -1)(5 +y)?